OpenStudy (anonymous):

Solve the system of equations. y = x + 5 y = x2 + 9x + 17

OpenStudy (nubeer):

subsitute the value of y in second equation means u have y = x+5 in first equation subsitute y in second one x+5 = x^2 + 9x +17 now find the value of x